Which of the following NaOH solutions have the highest pH value a) 1M b) 2M c) 0.5M d) 0.1
Answer #1

Answer

b) 2M

Explanation

NaOH is a strong base and completely dissociate ,so

[NaOH] = [OH-]

pOH = -log[OH-]

pH = 14 - pOH

as OH- concentration increases pH increases

a) pOH = - log(1M)

pOH = 0

pH = 14 - 0 = 14

b) pOH = - log(2M)

pOH = - 0.3010

pH = 14 - (-0.3010) = 14.30

c) pOH = -log(0.5M)

pOH = 0.30

pH = 14 - 0.30 = 13.7

d) pOH = - log( 0.1M)

pOH = 1

pH = 14 - 1 = 13

Therefore ,

The answer is b) 2M